The sum of the squares of three consecutive positive even integers is 116.
What are the numbers?

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

The sum of the squares of three consecutive positive even integers is 116.
What are the numbers?
------------------
116/3 = apx 39
sqrt(39) = apx 6, the middle number
--> 4, 6 & 8

Answer by ewatrrr(24785)


Hi
Let x,(x+2),(x+4)represent the three consecutive positive even integers
Question states***
x^2 + (x+2)^2 + (x+4)^2 = 116
Solving for x
3x^2 + 12x + 20 = 116
3x^2 + 12x - 96 = 0
x^2 + 4x - 32 = 0
factoring
(x + 8)(x-4)=0 Note:SUM of the inner product(8x) and the outer product(-4x) = 4x
(x + 8)=0 x = -8 is an Extraneous solution (not 'positive')
(x-4)=0 x = 4 the three consecutive positive even integers are 4,6,8
CHECKING our Answer***
16 + 36 + 64 = 116